#1.


A 2.0 kg particle moves as a function of time as follows:


x = 3.5 sin (1.5t+π/2)


Where, distance is measured in meters and time in seconds.


a)What are the equations for the potential and kinetic energies of the particle in Question #1?






1
Expert's answer
2021-05-11T16:35:08-0400
"K=0.5mv^2\\K=0.5(2)(-5.25\sin{1.5t})^2=27.5625\sin^2{1.5t}"

"U=0.5kx^2=27.5625\cos^2{1.5t}"